# Video Soundtrack Generation by Aligning Emotions and Temporal Boundaries

### v2

@@ 3.1 Conditional music generator
Our model generates music that is synchronized with the video’s temporal boundaries. We define temporal boundaries as scene cut locations in video and long-duration chord locations in music. Since we employ a hybrid approach, we first train a music generator capable of incorporating boundary locations as input and producing music with chords near these boundaries. To achieve this, we label the chords in the Lakh Pianoroll Dataset. During training, we consider only guitar and piano chords with at least three simultaneous notes that last for a minimum of two beats. These chords are labeled by inserting a CHORD token before the first ON token of each chord. To allow the model to generate chords not only at video scene cut locations but also independently where musically appropriate, we randomly remove 20% of CHORD tokens during training. An example of a labeled chord is shown in the top row of Figure 3. In training, the temporal locations of chords from the ground-truth MIDI serve as input boundaries, whereas during video-based inference, we replace these boundaries with video scene cut locations.
@@
Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following (Bharucha and Krumhansl 1983). During inference, forcing a CHORD token into the sequence at a specific location may cause the chord to sound off-beat or overly abrupt. This occurs because the model, having no prior knowledge of the upcoming chord, may generate preceding notes that do not align naturally with it. To address this, we develop a method that enables the model to “anticipate” upcoming chords and generate preceding notes and time shifts accordingly. Additionally, we train the model to generate the CHORD token itself, ensuring rhythmic consistency in the generated music. To achieve this, we define boundary offsets for each input token, representing the remaining time until the next boundary. These offsets are capped at a maximum value and, since our music generator is autoregressive, they are computed based on future chords rather than past ones. Furthermore, we amplify the loss of the CHORD token by a factor of 10, as it influences multiple preceding chord notes and plays a critical role in structuring the generated music. Sample offsets are illustrated in the bottom row of Figure 3.
@@
We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings (Wang et al. 2020) along the feature dimension, as shown in the lower part of Figure 2. This architectural choice is motivated by several key factors. First, similar to the emotion-conditioning mechanism, we inject boundary offset encodings at the input level rather than within the transformer body, ensuring that the core model remains unchanged. This allows the model to process inputs with or without boundary offsets, enabling seamless fine-tuning by repeating the learned positional encodings along the feature dimension when boundary offsets are absent. Second, we avoid adding boundary offset encodings directly to learned positional encodings to maintain a distinction between the two. Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them (Haviv et al. 2022; Kazemnejad et al. 2024). Based on this insight, we halve the feature length of learned positional encodings and allocate the remaining feature space to boundary offset encodings. The resulting vector sequence consists of positional encodings augmented with boundary offset encodings. Following the standard transformer model, we add this sequence to the token embeddings (Vaswani et al. 2017) before passing it to the transformer body with relative global attention (Huang et al. 2018). The transformer’s input is represented as:
$\begin{split}\mathbf{X}_{\text{trans}}&=\text{Concat}_{s}\left(\text{FFN}_{v}(x_{v}),\text{FFN}_{a}(x_{a}),\text{Embedding}(\mathbf{x}_{t})\right)\\ &+\text{Concat}_{f}(\text{FFN}_{b}(\mathbf{b})+\mathbf{W_{pe}})\end{split}$
@@
where $x_{v}$ and $x_{a}$ are the valence and arousal inputs, $\mathbf{x}_{t}$ is the input token sequence, $\mathbf{b}$ represents the input boundary offsets, and $\mathbf{W_{pe}}$ is the learned positional encoding. The feed-forward networks $\text{FFN}_{v}$, $\text{FFN}_{a}$, and $\text{FFN}_{b}$ process the valence, arousal, and boundary offset inputs, respectively. The operations $\text{Concat}_{s}$ and $\text{Concat}_{f}$ denote concatenation along the sequence and feature dimensions, respectively.
@@
Figure 3: Graphical illustration of boundaries. Top: In symbolic music, a chord with three or more simultaneous notes and a duration exceeding a set threshold defines a musical boundary and is used during training. Middle: In video, scene cuts serve as video boundaries and are used during inference. Bottom: Boundary offset represents the temporal distance to the next boundary. These figures are illustrative; offsets do not perfectly align with the music except at the boundary.
@@
Algorithm 1 outlines the process of computing boundary offsets during inference, i.e., music generation. The model generates music autoregressively, producing one token per forward pass. We maintain a time cursor by tracking the generated TIMESHIFT tokens and compute the boundary offset, i.e., the time remaining until the next boundary, for each generated token. If the model generates a CHORD token, we calculate the absolute difference between the time cursor and each input boundary. The boundary with a difference smaller than the predefined sensitivity threshold of 1 s is considered successfully generated. We then remove these boundaries from future offset calculations by replacing them with infinity. Next, we compute the boundary offset for the generated token, regardless of its type. This offset is determined as the distance to the next closest boundary, capped by a maximum offset value. For simplicity, Algorithm 1 is presented for a single sample, but in practice, this operation is performed in minibatches. During training, we preprocess the entire input sequence at once by constructing a time grid instead of a time cursor, allowing us to calculate boundary offsets for all tokens simultaneously. For clarity, the initial list of generated tokens is shown as empty; however, in practice, we begin with a START token.
@@
The ON tokens that appear after a CHORD token and before the next TIMESHIFT token are considered notes of the generated chord. To make the generated chord more distinctive, we increase the velocity of these notes. Music generation continues until the time cursor reaches the duration of the video. After synthesizing the generated MIDI into an audio waveform, we apply a 3-second fade-out effect to prevent an abrupt ending and ensure the music matches the duration of the input video.
@@
Parameter: Sensitivity, $\xi$; maximum offset, $\delta_{max}$; generated boundary mask, $\mathbf{m_{b}}$; token type $t_{t}$; token value, $t_{v}$; time cursor, $c$; generation function including forward-pass and sampling, g
@@
   $\mathbf{m_{b}}=\left|t-\mathbf{b}\right|<\xi$.
@@
3.2 Scene detection
During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations. As scene detection is a solved problem, we employ the off-the-shelf FFmpeg software [^1] https://www.ffmpeg.org/. As we generate musical chords near scene boundaries, we aim to avoid an excessive number of densely spaced chords. We therefore apply a difference filter to the extracted scene cuts and remove those occurring less than 4 s apart. The remaining timestamps serve as the input boundaries for the music generator.
3.3 Emotion mapping
We use a pretrained emotion classifier to extract emotions from input videos as probability distributions over discrete emotion categories (Sulun, Viana, and Davies 2024). Our music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories. Since the emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane (Russell 1980) before feeding them into the music generator.
Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories (1977). They presented their findings in a table that contains the mean and standard deviation values of valence and arousal for each categorical emotion. We extracted and used the values corresponding to Ekman’s six basic emotions (1971), which are presented in Table 1.
@@ 3.3 Emotion mapping
Table 1: Means and standard deviations of valence and arousal values corresponding to Ekman’s categorical emotions (1971), obtained from the user studies of Russell and Mehrabian (1977).
Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions. Valence-arousal values can be obtained either by sampling an emotion category based on the output probabilities and then drawing from its corresponding Gaussian, or by using the mixture mean computed as the weighted average of the category means. Additionally, we introduce a parameter that represents the maximum absolute value of means across all emotion categories. Using this parameter, we compute a single scaling coefficient applied to all valence-arousal distributions. This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
Although not used in our work, it is possible to invert this mapping, that is, to map valence-arousal values to emotion categories by comparing the valence-arousal coordinates to each category’s distribution defined by its mean and standard deviation. This involves calculating a distance metric, such as Euclidean or Mahalanobis distance, or computing the likelihood assuming Gaussian distributions, and then assigning the category with the smallest distance or highest likelihood.
4 Experimental setup
@@
We perform objective and subjective evaluations of our overall model using two datasets, against the open-source state-of-the-art models that generate MIDI from arbitrary videos: Video2Music (Kang, Poria, and Herremans 2024) and CMT (Di et al. 2021).
4.1 Implementation details
Our music generator consists of 11 layers, 8 attention heads, and a model dimensionality of 512. To ensure a fair comparison, we align our model size with the compared models: our model has 37M parameters, compared to 39M in CMT (Di et al. 2021) and 33M in Video2Music (Kang, Poria, and Herremans 2024). Training is done with a context length of 1216 and a batch size of 64. The model is implemented in PyTorch (Paszke et al. 2019) and trained on a single NVIDIA A100 80GB GPU.
For video-based inference in evaluation, we set the maximum absolute value of the emotion category means to 0.8. To ensure consistency, all methods generate MIDI outputs that are synthesized into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com, followed by peak normalization to -3 dB. The audio is then merged with the input video by overwriting its audio channel. For all methods, outputs are generated in a single run across all datasets using a fixed random seed.
@@ 4.1 Implementation details
We train our model using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps. We use the Adam optimizer with gradient clipping to a norm of 1 (Kingma and Ba 2015). For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen number of semitones between -3 and 3, inclusive. We reserve 5% of the training data for validation. Hyperparameters are selected via grid search to minimize validation loss.
4.2 Evaluation datasets
@@
For objective and subjective evaluation, we first use the EmoMV-C dataset, which consists of music videos containing music audio tracks to be used as ground truth for objective evaluation (Thao, Roig, and Herremans 2023). We exclude auditory features during the emotion classification to prevent data leakage. To assess the model’s generalization across different video types, we also include the Pittsburgh Advertisements (Ads) dataset, which contains advertisement videos that may or may not feature music (Hussain et al. 2017). As advertisements often rely on music to maximize viewer engagement, the Ads dataset provides a real-world scenario for evaluation. (Zander 2006).
@@
We use the full EmoMV-C validation split, which contains 48 thirty-second videos. We filter the Ads dataset to match the total duration of EmoMV-C. We select videos associated with the four basic emotions commonly used in music emotion classification: cheerful, calm, angry, and sad (Yang et al. 2008). We then remove videos shorter than 1 minute. To ensure an unbiased selection, we use YouTube IDs, which are generated randomly by YouTube. We sort these IDs alphabetically and select the first six videos from each emotion category, resulting in 24 evaluation videos. Finally, we trim the videos to a uniform duration of 1 minute. The resulting evaluation videos are then fed into the compared models.
4.3 Objective evaluation

### v3

@@ III-B2 Temporal boundary matching
Our model generates music that is synchronized with the video’s temporal boundaries. We define temporal boundaries as scene cut locations in video and long-duration chord locations in music. Since we employ a hybrid approach, we first train a music generator capable of incorporating boundary locations as input and producing music with chords near these boundaries. To achieve this, we label the chords in the Lakh Pianoroll Dataset. During training, we consider only guitar and piano chords with at least three simultaneous notes that last for a minimum of two beats. These chords are labeled by inserting a CHORD token before the first ON token of each chord. An example of a labeled chord is shown in the top row of Figure 3. In training, the temporal locations of chords from the ground-truth MIDI serve as input boundaries, whereas during video-based inference, we replace these boundaries with video scene cut locations.
@@
Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following [2]. During inference, forcing a CHORD token into the sequence at a specific location may cause the chord to sound off-beat or overly abrupt. This occurs because the model, having no prior knowledge of the upcoming chord, may generate preceding notes that do not align naturally with it. To address this, we develop a method that enables the model to “anticipate” upcoming chords and generate preceding notes and time shifts accordingly. Additionally, we train the model to generate the CHORD token itself, ensuring rhythmic consistency in the generated music. To achieve this, we define boundary offsets for each input token, representing the remaining time until the next boundary. These offsets are capped at a maximum value and, since our music generator is autoregressive, they are computed based on future chords rather than past ones. Sample offsets are illustrated in the bottom row of Figure 3.
@@
We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings [49] along the feature dimension, as shown in the lower part of Figure 1. This architectural choice is motivated by several key factors. First, similar to the emotion-conditioning mechanism, we inject boundary offset encodings at the input level rather than within the transformer body, ensuring that the core model remains unchanged. This allows the model to process inputs with or without boundary offsets, enabling seamless fine-tuning by repeating the learned positional encodings along the feature dimension when boundary offsets are absent. Second, we avoid adding boundary offset encodings directly to learned positional encodings to maintain a distinction between the two. Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them [13, 22]. Based on this insight, we halve the feature length of learned positional encodings and allocate the remaining feature space to boundary offset encodings. The resulting vector sequence consists of positional encodings augmented with boundary offset encodings. Following the standard transformer model, we add this sequence to the token embeddings [48] before passing it to the transformer body with relative global attention [18]. The transformer’s input is represented as:
$\begin{split}\mathbf{X}_{\text{trans}}&=[\text{FFN}_{v}(x_{v}),\,\text{FFN}_{a}(x_{a}),\,\text{Embedding}(\mathbf{x}_{t})]_{1}\\ &\quad+[\text{FFN}_{b}(\mathbf{b}),\,\mathbf{W}_{\text{pe}}]_{2}\end{split}$
$\text{FFN}_{v}(x_{v}),\text{FFN}_{a}(x_{a})\in\mathbb{R}^{1\times D},\quad\text{Embedding}(\mathbf{x}_{t})\in\mathbb{R}^{L\times D},$
$\text{FFN}_{b}(\mathbf{b}),\mathbf{W}_{\text{pe}}\in\mathbb{R}^{(L+2)\times\frac{D}{2}},\quad\;\;\,\mathbf{X}_{\text{trans}}\in\mathbb{R}^{(L+2)\times D}$
@@
where $x_{v}$ and $x_{a}$ are the valence and arousal inputs, $\mathbf{x}_{t}$ is the input token sequence, $\mathbf{b}$ represents the boundary offsets, and $\mathbf{W}_{\text{pe}}$ is the learned positional encoding. The feed-forward networks $\text{FFN}_{v}$, $\text{FFN}_{a}$, and $\text{FFN}_{b}$ process the valence, arousal, and boundary offset inputs, respectively. The plus sign ++ denotes tensor addition. The first concatenation $\cdot\,]_{1}$ is along the sequence (first) dimension producing a sequence of length $L+2$, and the second concatenation $\cdot\,]_{2}$ is along the feature (second) dimension producing a feature of size $D$.
@@
Figure 3: Illustration of boundaries. Top: In symbolic music, a chord with three or more simultaneous notes and a duration exceeding a set threshold defines a musical boundary and is used during training. As CHORD is a special token without duration or pitch, it is omitted from the pianoroll representation. Its temporal location, however, aligns with the red line indicating the boundary. Middle: In video, scene cuts serve as video boundaries and are used during inference. Bottom: Boundary offset represents the temporal distance to the next boundary. Each offset is a scalar value, visualized as the height of the corresponding green bar. These figures are illustrative; offsets do not perfectly align with the music except at the boundary.
@@
Algorithm 1 outlines the process of computing boundary offsets during inference, i.e., music generation. The model generates music autoregressively, producing one token per forward pass. We maintain a time cursor by tracking the generated TIMESHIFT tokens and compute the boundary offset, i.e., the time remaining until the next boundary, for each generated token. If the model generates a CHORD token, we calculate the absolute difference between the time cursor and each input boundary. The boundary with a difference smaller than the predefined sensitivity threshold $\xi$ is considered successfully generated. We then remove these boundaries from future offset calculations by replacing them with infinity. Next, we compute the boundary offset for the generated token, regardless of its type. This offset represents the distance to the next closest boundary. We compute it by subtracting the time cursor from each input boundary and replacing any negative values (corresponding to past boundaries) with infinity to ignore them. We then take the minimum of the resulting values as the offset to the nearest boundary and clamp it to a maximum value if necessary. The resulting boundary offset and generated token are appended to their respective lists to be used in the next timestep. For simplicity, Algorithm 1 is presented for a single sample, but in practice, this operation is performed in minibatches. During training, we preprocess the entire input sequence at once by constructing a time grid instead of a time cursor, allowing us to calculate boundary offsets for all tokens simultaneously. For clarity, the initial list of generated tokens is shown as empty; however, in practice, we begin with a START token.
@@
The ON tokens that appear after a CHORD token and before the next TIMESHIFT token are considered notes of the generated chord. To make the generated chord more distinctive, we increase the velocity of these notes. Music generation continues until the time cursor reaches the duration of the video.
@@
Parameter: Sensitivity, $\xi$; distances to all boundaries, $\boldsymbol{\delta}_{b}$; maximum offset, $\delta_{\max}$; generated boundary mask, $\mathbf{m_{b}}$; token type $t_{t}$; token value, $t_{v}$; time cursor, $c$; generation function including forward-pass and sampling, g
@@
   $\mathbf{m_{b}}=\left|c-\mathbf{b}\right|<\xi$.
@@
  $\boldsymbol{\delta}_{b}=\mathbf{b}-c$
  $\boldsymbol{\delta}_{b}[\boldsymbol{\delta}_{b}<0]=+\infty$
@@
III-B3 Scene detection
During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations. As scene detection is a well-studied problem with established solutions, we employ the off-the-shelf FFmpeg software [^1] https://www.ffmpeg.org/. As we generate musical chords near scene boundaries, we aim to avoid an excessive number of densely spaced chords. We therefore apply a difference filter to the extracted scene cuts and remove those occurring less than $t_{\text{diff}}$ seconds apart. The remaining timestamps serve as the input boundaries for the music generator.
III-B4 Emotion mapping
We use our pretrained emotion classifier [44] to extract emotions from input videos as probability distributions over Ekman’s basic emotion categories [10]. Our music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories. Since the emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane [38] before feeding them into the music generator.
Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories [37]. They presented their findings in a table that contains the mean and standard deviation values of valence and arousal for each categorical emotion. We extracted and used the values corresponding to Ekman’s six basic emotions [10], which are presented in Table I. While Russell and Mehrabian presented one of the oldest user studies that examine the relationship between dimensional and categorical emotion representations, their study remains the most comprehensive in terms of number of users ($N=300$) and number of categorical emotions ($K=151$), and is still employed in recent machine learning applications [27]. Similar and newer user studies from Hoffmann et al. ($N=70$, $K=22$) [15] and Trnka et al. ($N=187$, $K=16$) [47] are more limited, and do not include all of Ekman’s basic emotions.
Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions. Valence-arousal values can be obtained either by sampling an emotion category based on the output probabilities and then drawing from its corresponding Gaussian, or by using the mixture mean computed as the weighted average of the category means. Additionally, we introduce a parameter $\mu_{\max}$ representing the maximum absolute value of means across all emotion categories. Using $\mu_{\max}$, we compute a single scaling coefficient applied uniformly to all valence-arousal distributions. This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
Although not used in our model, we also formalize a method to invert this mapping, that is, to map valence-arousal values to emotion categories by comparing the valence-arousal coordinates to each category’s distribution defined by its mean and standard deviation. This involves calculating a distance metric, such as Euclidean or Mahalanobis distance, or computing the likelihood assuming Gaussian distributions, and then assigning the category with the smallest distance or highest likelihood. This provides a generalizable framework for integrating dimensional and categorical emotion representations in multimodal systems.
TABLE I: Valence and arousal means and standard deviations for Ekman’s six basic emotions [10], derived from Russell and Mehrabian’s user study [37].
@@ III-B4 Emotion mapping
IV Experimental setup
@@
We perform objective and subjective evaluations of our overall model using two datasets, against open-source state-of-the-art models that generate MIDI from arbitrary videos. Comparison with V-MusProd [56] and XMusic [46] was not possible, as their code is not publicly available. We therefore compare our method against Video2Music [21] and CMT [8].
IV-A Implementation details
@@
Our music generator consists of 11 layers, 8 attention heads, and a model dimensionality of 512. To ensure a fair comparison, we align our model size with the compared models: our model has 37M parameters, compared to 39M in CMT [8] and 33M in Video2Music [21]. Training is done with a context length of 1216 and a batch size of 64. The model is implemented in PyTorch [33] and trained on a single NVIDIA A100 80GB GPU. We train our model using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps. We use the Adam optimizer with gradient clipping to a norm of 1 [24]. For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen number of semitones between -3 and 3, inclusive. To prevent the model from becoming overly reliant on the CHORD token for chord generation, we randomly remove 20% of these tokens during training. Furthermore, we amplify the loss of the CHORD token by a factor of 10, as it influences multiple preceding chord notes and plays a critical role in structuring the generated music. We set the maximum boundary offset $\delta_{\max}$ to 8 seconds. We reserve 5% of the training data for validation. These training hyperparameters are selected via grid search to minimize validation loss.
For video-based inference in evaluation, we set the maximum absolute value of the emotion category means $\mu_{\max}$ to 0.8. We obtain valence–arousal values as the mean of the weighted mixture of Gaussians derived from the predicted emotion categories. Although this conditioning is deterministic, the music generator remains stochastic due to probabilistic token sampling during generation. The difference filter discards scene timestamps that are less than $t_{\text{diff}}=4$ seconds apart. We set the boundary sensitivity threshold $\xi$ to 1 second. To ensure consistency, all methods generate MIDI outputs that are synthesized into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com, followed by peak normalization to -3 dB. We apply a 3-second fade-out effect to prevent an abrupt ending and ensure the music matches the duration of the input video. The audio is then merged with the input video by overwriting its audio channel. For all methods, outputs are generated in a single run across all datasets using a fixed random seed. The inference hyperparameters are chosen by manually inspecting outputs on videos not included in the evaluation.
IV-B Evaluation datasets
@@
For objective and subjective evaluation, we first use the EmoMV-C dataset, which consists of music videos containing music audio tracks to be used as ground truth for objective evaluation [45]. We exclude auditory features during the emotion classification to prevent data leakage. To assess the model’s generalization across different video types, we also include the Pittsburgh Advertisements (Ads) dataset, which contains advertisement videos that may or may not feature music [20]. As advertisements often rely on music to maximize viewer engagement, the Ads dataset provides a real-world scenario for evaluation [54].
@@
We use the full EmoMV-C validation split, which contains 48 thirty-second videos. We filter the Ads dataset to match the total duration of EmoMV-C. We select videos associated with the four basic emotions commonly used in music emotion classification [52], covering the quadrants of Russell’s valence-arousal model [38]: cheerful, calm, angry, and sad, providing a broad coverage of the emotional space for evaluation. While some of the chosen emotions do overlap with Ekman’s categories [10], we avoid explicitly using Ekman’s emotions to select the evaluation videos. This prevents biasing the evaluation toward our method, since our video emotion classifier is trained to output Ekman’s categories. We exclude videos shorter than 1 minute. To ensure an unbiased selection, we use YouTube IDs, which are generated randomly by YouTube. We sort these IDs alphabetically and select the first six videos from each emotion category, resulting in 24 evaluation videos. Finally, we trim the videos to a uniform duration of 1 minute. The resulting evaluation videos are then fed into the compared models.
Figure 4: Sample view of the survey used in the subjective evaluation. Participants were asked to drag the method boxes from the left onto the ranking boxes on the right.
IV-C Objective evaluation